OpenStudy (anonymous):

Marta has $6000 to invest. She puts x dollars into an account that earns 2% interest each year. She puts the rest into a COD (another account) that earns 4% interest each year. Suppose she has a total of $6210 at the end of the year. How much did she invest in each account? Please show work.

OpenStudy (anonymous):

2%x + 4% ( 6,000 - x) = $210

OpenStudy (anonymous):

Can you continue with it?

OpenStudy (anonymous):

Yeah, thanks

OpenStudy (anonymous):

:)

OpenStudy (anonymous):

(:

OpenStudy (anonymous):

You'll have Harkirat solution to double check your work, happy :)

OpenStudy (anonymous):

@Harkirat thanks you!

OpenStudy (anonymous):

Let her invest $x in the first account then SI of x will be = PRT = x*0.02*1 = 0.02x amount = x + 0.02x = 1.02x Then sum invested in second account = $(6000 -x) so SI = PRT = (6000-x)*0.04*1 = 240 - 0.04x amount = (6000 - x) + (240 - 0.04x) = 6240 - 1.04x ATQ 1.02x + (6240 - 1.04x) = 6210 1.02x - 1.04x = 6210 - 6240 -0.02x = -30 x = -30/-0.02 x = 1500 so she invested 1500 in the first account giving 2% interest and 4500 in the second account giving 4% interest

OpenStudy (anonymous):

Thanks, I just got that(:

OpenStudy (anonymous):

good...☺ n thanks..